OpenStudy (australia10):

How do we turn 0.151 repeating into a decimal ? ​ ​​

OpenStudy (danjs):

you mena a fraction?

OpenStudy (jackthegreatest):

you cant, unless if you had the faction

OpenStudy (jackthegreatest):

oh or do you mean 0.151151151

OpenStudy (australia10):

Fraction * sorry

OpenStudy (jackthegreatest):

ahh haha

OpenStudy (jackthegreatest):

and your done :D

OpenStudy (jackthegreatest):

because it is .151, which is in the thousands place, thus dividing by a thousand

OpenStudy (australia10):

how does khun get 151/999

OpenStudy (jackthegreatest):

oh whoops repeating decimal

OpenStudy (danjs):

Let s be the repeating decimal.. S=0.151... 1000*s = 151.151... If you subtreact 1000*s - s, the repeating decimal is eliminated \[1000*s - s = 151.151151151... - 0.151151151....\]

OpenStudy (danjs):

\[999*s = 151\] s = 151 / 999

OpenStudy (danjs):

similar process for all the other repeating decimals, multiply it by 10, 100,1000, or whatever to get one whole number of that repeating part

OpenStudy (australia10):

Thsanks